OpenStudy (anonymous):

solve...simultaneous equations 3x+4y=16 6x+15y=-6

OpenStudy (anonymous):

i already got to the answer -11y=22 wat happens now??

OpenStudy (anonymous):

Well, you said you got the answer, but let me do it over, and see if you are right.:)

OpenStudy (anonymous):

ta!!! but is till need to find out wat x is

OpenStudy (anonymous):

Oh, I see, okay, let me do this step by step. Ok, first we have to get either the x's or y's to cancel out, so lets multiply the first equation by 2 to cancel out the x's.

OpenStudy (anonymous):

oh sorry i meant wat y is

OpenStudy (anonymous):

3x + 4y = 16 x 2 x 2 x2 Ok, so we have now: 6x + 8y = 32 6x + 15y = -6 Ok, now what does the equations look like after canceling out the x's?

OpenStudy (anonymous):

@chayareisner And by canceling out the x's, we make it possible to solve for y.

OpenStudy (anonymous):

thnx but ive already got -11y=22 i need to find one y

OpenStudy (anonymous):

Ok, well from there, you divide both sides by -11 to find y. -11y = 22 /-11 /-11

OpenStudy (anonymous):

rite so 22 share by -11 is it -2??

OpenStudy (anonymous):

Yup! y = -2

OpenStudy (anonymous):

Now substitute y for -2 to solve for x.

OpenStudy (anonymous):

ye ill do that alone thnx!!!